Please I need help this one ASAP

6 "A company bought a machine for $100,000. The machine was depreciated using a 5 year MACRS approach. After 4 years, the machine was sold at a salvage value of $91,300. Assuming a tax rate of 21%, what are the net proceeds from the sale of the machine?
(Hint: You will want to take your salvage values and add/subtract gains due to the sale.)"
Answer #1

book value at the end of year 4

= cost - accumulated depreciation

= 100000 - 100000 * (0.2+0.32+0.1920+0.1152)

= 17280

Tax on sale = (91300 - 17280) * 0.21

= 15544.2

net proceeds from sale

= 91300 - 15544.2

= 75755.8